QUESTION ONE

Mrs.Chausiku Mwendapole purchased share of ALI BABA for TZS 700 and received 

a dividend per share of TZS 100 on 31st December 2020. Currently the share price is 

worth TZS 800.

REQUIRED

Calculate the returns in TZS and percentage from the investment


1
Expert's answer
2021-03-26T08:08:02-0400

"P1 = 700"; "P2 = 800"; "R = 100"

Let's find the percentage change "\Delta"

"% = (P2-P1)\/P1""\Delta = (P2-P1)/P1 = 100/700= 0.1428"

Let's the returns "ROI = R\/P1 = 100\/700 = 0.1428"

Answer: both ROI and the percentage change are 14.3%